OpenStudy (anonymous):

How high is a wall that casts a 9 feet shadow at the same time that 4 feet flag pole cast a 4 feet shadow HHHHHEEEEELPPPPPPPPP

OpenStudy (anonymous):

x/9=4/4

OpenStudy (anonymous):

x/9=1 = 9 i think :P

OpenStudy (anonymous):

Yhu were right...awww thanks

OpenStudy (anonymous):

\[tg \theta a = tg \theta b \rightarrow x/9=4/4\]